OpenStudy (angela210793):

I need some help cause i have found a different no. not the 1 my classmates did...:S The arithmetic average of a and b is 9,the arithmetic average of a and c is 7.what is the arithmetic average of a,b and c?

OpenStudy (anonymous):

a + b = 18 a + c = 14

OpenStudy (angela210793):

Hmm it's not what i am looking for....i am looking for (a+b+c)/3...arithmetic average of a,b and c..thnx anyway :)

OpenStudy (anonymous):

i know but this is a starting point

OpenStudy (angela210793):

ooh...right :) sorry :)

myininaya (myininaya):

do we not have another equation? i only see two equations and 3 unknowns we need another equation if you a number that is known like because we can write the answer right now as (18+c)/3 or (32-a)/3 we can also write it in terms of b

OpenStudy (angela210793):

no that's what we have...

myininaya (myininaya):

so do you want the answer in terms of c, a, or b?

OpenStudy (angela210793):

none of that...cause we have to find a integer....

OpenStudy (anonymous):

23/3

myininaya (myininaya):

how do you get c=5, anwar?

OpenStudy (angela210793):

Anwar i's appreciate it if u'd explain that please.....

myininaya (myininaya):

or a=9

myininaya (myininaya):

or b=9

OpenStudy (angela210793):

OMG!!!!!

OpenStudy (anonymous):

Well, since we will add them all up after all, we can choose any values if a,b and c that satisfy the two given equation \(a+b=18\) and \(a+c=14\). So choosing \(a=b=9\) would imply \(c=5\), then \((a+b+c)/3=23/3\).

OpenStudy (angela210793):

Will u please forgive me...my classmate just e-mailed me and i have forgotten smth....b+c=8...soo sorry

OpenStudy (anonymous):

This does not seem true :(

OpenStudy (anonymous):

thats it, so myininaya was right :)

OpenStudy (angela210793):

yea....

OpenStudy (angela210793):

I think i can solve it by my own now...:) Thanks a lot..i appreciate it :):):)

OpenStudy (anonymous):

ans anwarA u r just guessing the values, a+b=18. even if a=8 b=10

OpenStudy (angela210793):

i found it....it's 8

myininaya (myininaya):

22/3

OpenStudy (anonymous):

What I said was totally wrong and stupid. :(

myininaya (myininaya):

oh wait i made a mistake

OpenStudy (angela210793):

(a+b)/2+(b+c)/2+(a+c)/2=9+8+7=24 2(a+b+c)/2=24 (a+b+c)/3=24/3=8 :)

OpenStudy (angela210793):

this is how u do it :) thanks anyway :)

OpenStudy (anonymous):

Oh you didn't give that \(b+c=8\) -.-

OpenStudy (anonymous):

\((b+c)/2=8\)*

OpenStudy (angela210793):

(b+c)/2=8...didn't i wrote this way?

OpenStudy (anonymous):

a+b=18 a+c=14 ---------- b-c=4 b+c=8 -------- b=6 c=2 a=12 6+2+12/3

myininaya (myininaya):

is the average of the two 8 or is b+c=8

OpenStudy (anonymous):

How can we get b+c=8? Please explain

OpenStudy (angela210793):

Guys..I'm soo sorry...it's morning....haven't had breakfast yet...so....it seems my brain isn't working well...:P So sorry...:(:(:(:(:(

OpenStudy (angela210793):

the average of b+c=8

OpenStudy (angela210793):

Going to eat...thnx :)

myininaya (myininaya):

ok so 24/3=8 is correct gj because we get a=8 b=10 c=6 so (8+10+6)/3=8

OpenStudy (anonymous):

a+b=18, a+c=14 and b+c=16 Hence, 2a+2b+2c=18+14+16=48 Or, a+b+c=24 Therefore, (a+b+c)=24/3=8

OpenStudy (angela210793):

Thank you guys :)